Grade 11General Physics

Determine the volume contraction of a solid copper cube, 10 cm on an edge, when subjected to a hydraulic pressure of 7.0 × 106 Pa.

Jayant Kumar
12 Years agoGrade 11

1 Answer

arpit Patel
9 Years ago
Here we know that B=PV/∆v So, ∆v=PV/B Here B is equal to 140×10race to power 9 pa P=7×10 race to power 6 paL=10race to power -1 m so through L we can find the volume formula =cube of side I.e. 10 race to power -3So through first formula we can solve it But by solving it will =0.65×10race to power -6 I hope that I had helped you Thank you